OpenStudy (anonymous):

What is the exact value of sqrt 54/3 sqrt 2

jhonyy9 (jhonyy9):

54=2*27=2*3*9=2*3*3*3=2*3^3 so than sqrt54 will be equale with ?

jhonyy9 (jhonyy9):

can you calcule it ?

jhonyy9 (jhonyy9):

sqrt54 =sqrt(2*3*3^2) = ?

jhonyy9 (jhonyy9):

sqrt54 =3sqrt6

jhonyy9 (jhonyy9):

can you simplifie it now ?

jhonyy9 (jhonyy9):

3sqrt6 ------ = ? 3sqrt2